Let A and B be two events defined on a sample space S. If Pr(A)=0.8; Pr(A| B )=0.85 and Pr(A| B^ c )=0.75; determine the probability that neither of the two events occur.


1
Expert's answer
2022-04-04T16:15:39-0400

The probability that neither of these events occur is: "P(\bar{A}\cap\bar{B})", We assume that "B^c=\bar{B}." By definition, we have "P(A|\bar{B})=\frac{P(A\cap \bar{B})}{P(\bar{B})}". "P(A|{B})=\frac{P(A\cap {B})}{P({B})}". The following formula holds: "P(\bar{B})=P({A}\cap \bar{B})+P({\bar{A}}\cap {\bar{B}})". From the latter we get: "P(\bar{B})-P({A}\cap \bar{B})=P({\bar{A}}\cap {\bar{B}})". "P({A}\cap \bar{B})=P(A|\bar{B})P(\bar{B})". Thus, we receive that: "P({\bar{A}}\cap {\bar{B}})=P(\bar{B})(1-P(A|\bar{B}))=(1-P({B})(1-P(A|\bar{B}))"."P(A)=P({A}\cap \bar{B})+P({{A}}\cap {{B}})". From formulae for "P(A|\bar{B})" and "P(A|{B})" we receive: "P(A)=P(A|\bar{B})(1-P({B}))+P(A|{B})P(B)". From the latter we receive:"P(B)=\frac{(P(A)-P(A|\bar{B}))}{P(A|B)-P(A|\bar{B})}". Thus, we have: "P(B)=\frac{0.8-0.75}{0.85-0.75}=0.5". Substitute the latter into expression for "P({\bar{A}}\cap {\bar{B}})": "P({\bar{A}}\cap {\bar{B}})=0.5\cdot0.25=0.125." Thus, the probability, that neither of events occur is: "0.125."